OpenStudy (anonymous):

Six cans of soda and two bags of potato chips cost $7.08. And eight cans of soda and five bags of potato chips cost $12.45.What is the price per can for the soda, and price per bag of potato chips?

OpenStudy (anonymous):

6s + 2p = $ 7.08 ( 1) 8s + 5p = $12.45 ( 2) Multiply the first ( 1) for 5/2: 15s + 5p = 17.7 ( 3) ( 3) - ( 2): -> 7s = 5.25 => s = 5.25/ 7 = .75 Thus the cost per soda is 75 cents. Can you find the cost of potato?

OpenStudy (anonymous):

that confused me more....

OpenStudy (anonymous):

How? Why?